# DeWave: Discrete EEG Waves Encoding for Brain Dynamics to Text Translation

### v1

@@ Abstract
The translation of brain dynamics into natural language is pivotal for brain-computer interfaces (BCIs), a field that has seen substantial growth in recent years. With the swift advancement of large language models, such as ChatGPT, the need to bridge the gap between the brain and languages becomes increasingly pressing. Current methods, however, require eye-tracking fixations or event markers to segment brain dynamics into word-level features, which can restrict the practical application of these systems. These event markers may not be readily available or could be challenging to acquire during real-time inference, and the sequence of eye fixations may not align with the order of spoken words. To tackle these issues, we introduce a novel framework, DeWave, that integrates discrete encoding sequences into open-vocabulary EEG-to-text translation tasks. DeWave uses a quantized variational encoder to derive discrete codex encoding and align it with pre-trained language models. This discrete codex representation brings forth two advantages: 1) it alleviates the order mismatch between eye fixations and spoken words by introducing text-EEG contrastive alignment training, and 2) it minimizes the interference caused by individual differences in EEG waves through an invariant discrete codex. Our model surpasses the previous baseline (40.1 and 31.7) by $3.06\%$ and $6.34\%$, respectively, achieving 41.35 BLEU-1 and 33.71 Rouge-F on the ZuCo Dataset. Furthermore, this work is the first to facilitate the translation of entire EEG signal periods without needing word-level order markers (e.g., eye fixations), scoring 20.5 BLEU-1 and 29.5 Rouge-1 on the ZuCo Dataset, respectively. Codes and the final paper will be public soon.
Ground Truth
Bob attended the University of Texas at Austin where he graduated, Phi Beta Kappa with a Bachelor’s degree in Latin American Studies in 1973, taking only two and a half years to complete his work, and obtaining generally excellent grades.
Predict
Figure 1: Overall illustration of translate EEG waves into text through quantised encoding.
@@ 1 Introduction
Decoding brain states into comprehensible representations has long been a focal point of research [19, 37]. Electroencephalogram (EEG) signals are particularly favored by researchers due to their non-invasive nature and ease of recording [29, 45]. Traditional EEG decoding techniques largely focus on classifying brain states into restricted categories like Motor Imaginary (MI) [36, 7], Emotion [18, 46], Robotic Control [43], and Gaming [26, 22]. However, these labels, bound to specific tasks, are insufficient for broad-based brain-computer communication. Consequently, there has been a surge of interest in brain-to-text (speech) translation in recent years. As the current trend leans towards large language models (LLM) [4, 11, 28] showcasing increasingly generalized intelligence capabilities, it becomes crucial to delve into ways of bridging the gap between brain signals and natural language representation. However, this area remains under-explored.
@@ 2 Related Works
The key to decoding natural language from EEG signals is good representations. Existing work for EEG-to-text representation can be categorized into hand-crafted representations and deep-learning representations. Earlier works make use of traditional methods to obtain hand-crafted representations from a sequence of EEG signals. A continuous feature representation is extracted using methods such as statistical features [50], correlation coefficient matrix from fast Fourier transform (FFT) components [42], wavelet transform features [17, 30, 33], and Mel-frequency cepstral coefficients (MFCCs) [5]. These hand-crafted representations can be used to establish a mapping between EEG segments and words using distance-based methods such as linear discriminative analysis (LDA) or Support Vector Machine (SVM).
@@ 3.3 EEG Vectorization
Word-Level EEG Features With Event Markers:  The EEG waves are first sliced into fragments according to the eye-tracking fixation of word sequences given in the annotation. Similar to [47], we calculate the statistical result of four frequency band filters, Theta band (5-7Hz), the Alpha band (8-13Hz), the Beta band (12-30Hz), and Gamma band (30Hz-) [25] to get the statistic frequency features of each fragment. It is noted that although different fragments may have different wavelengths, the statistical results are the same (embedding size 840). A multi-head transformer layer is applied to project the embedding into feature sequences with latent size 512.
@@
For structure-wise, a conformer-based multi-layer encoder with specially designed hyperparameters is employed. The one-dimensional convolution layer processes the EEG waves to generate the embedding sequence 333perception field is roughly 200ms with overlap 100ms for each embedding, fusing the EEG channels into a unique embedding for each period. We apply bi-directional transformer attention layers to the sequence to capture temporal relations.
@@
To achieve a semantically meaningful codex, a cross-modality contrastive learning approach is introduced. Similar to CLIP [34, 21] between vision and language, it connects the EEG-tokenized codex embeddings and the text embeddings in a joint embedding space. To establish this alignment without relation annotation, each corresponding pair of EEG codex and text word2vec encoding is treated as a positive pair, while non-corresponding pairs are treated as negative. The model then learns to differentiate between these pairs. The distance between embeddings of positive pairs is minimized while that of negative pairs is maximized, as defined in Equation 4. Let $z_{q}$ denote the EEG embedding and $z_{t}$ denote the text embedding. For a given EEG-text pair $(i,j)$, the loss is defined in Equation 4
$L_{\rm{contrast}}=-\frac{1}{n}\sum log\left[\frac{\exp(\mathbf{s}_{ii}/\tau)}{\sum_{k=1}^{2N}\exp(\mathbf{s}_{ik}/\tau)}\right],\mathbf{s}_{ij}=\mathbf{z}_{q}(\mathbf{x}^{i})^{T}\mathbf{z}_{t}(j)$ (4)
@@ 4.2 Implementation Details
The code could be accessed through an anonymous link  [^5] https://anonymous.4open.science/r/DeWave-EEG-Translation-6C2E. For word-level EEG features, we use the $56$ tokens each with an $840$ embedding size. For raw EEG waves, we clip or pad the EEG waves up to sample point 5500 with a constant value of zero. A transformer layer with head number 8 and a $1\times 1$ convolutional layer are combined to fuse multiple EEG channels into an embedding sequence with size 512. DeWave uses a codex with size 2048 where each codex latent is an embedding with size 512. The ablation study (Section 4.6) gives a discussion about the codex size. All models are trained on Nvidia V100 and A100 GPUs. For the self-supervised decoding for raw waves, we use a learning rate of 5e-4 and a VQ coefficient of 0.25 for training 35 epochs. For training the codex (stage 1), DeWave uses a learning rate of 5e-4 for 35 epochs. For finetuning the translation (stage 2), DeWave uses a learning rate of 5e-6 for 30 epochs. We use the SGD as the optimizer for training all the models. Due to limited space, refer to Appendix B for more details.
@@ 4.3 Evaluation Metrics
We evaluate translation performance using NLP metrics, BLEU and ROUGE, as shown in Table 1. For word-level EEG features, we compare our results to EEG-to-Text [47], maintaining a consistent language model for fairness. In the absence of methods for raw EEG waves, we establish a ${}^{\dagger}$—by segmenting the entire EEG waves into a sequence embedding using a 200ms time window with a 100ms overlap. We adapt Wave2Vec, originally developed for speech recognition, to brain waves and compare it with our approach, DeWave. Furthermore, we adapt unsupervised raw EEG waves classification methods BENDR [10] and SCL [9], using SSL pre-training and feature extraction for comparison, underscoring the impact of discrete encoding.
Table 1: Evaluation metrics of EEG-to-Text translation under both word-level features input and raw waves input.
@@
Raw Waves:   DeWave represents an unprecedented endeavor in translating raw EEG waves directly, obviating the need for event markers. This is achieved through the application of a learned Wave2Vec model, which converts waves into discrete codex tokens. When benchmarked against the baseline method - which merely slices the wave according to a time window - DeWave exhibits marked improvement. It attains BLEU-$1,2,3,4$ scores of $20.51,10.18,5.16$, and $2.56$, respectively, surpassing the baseline by margins of $7.44~{}(+56.92\%)$, $4.44~{}(+76.1\%)$, $2.61~{}(+102.35\%)$, and $1.42~{}(+129.09\%)$. This underscores the superiority of a learned discrete embedding representation over the rudimentary time-window baseline. In comparison with cutting-edge self-supervised learning (SSL) methods, DeWave consistently outperforms them, including contrastive learning (CL)-based methods such as SCL [9], the original Wave2Vec [2], and BENDR [10]. Interestingly, DeWave does not simply compete with SSL methods but also enhances them. The performance can be further bolstered by integrating DeWave with CL training, thereby verifying the efficacy of the proposed discrete coding. This strategy will be adopted in our final version to further consolidate our results.
@@ 4.4 Cross-Subject Performance
Cross-subject performance is of vital importance for practical usage. To further report the We further provide a comparison with both baseline methods and a representative meta-learning (DA/DG) method, MAML [8], which is widely used in cross-subject problems in EEG classification below.
Table 2: Cross-subject performance average decreasing comparison on 18 human subjects, where MAML denotes the method with MAML training. The metric is the lower the better.
Calib Data Method Eye fixation $-\Delta(\%)$downarrow Raw EEG waves $-\Delta(\%)$downarrow
B-2 B-4 R-P R-F B-2 B-4 R-P R-F
$\times$ Baseline 3.38 2.08 2.14 2.80 7.94 5.38 6.02 5.89
$\checkmark$ Baseline+MAML [8] 2.51 1.43 1.08 1.23 6.86 4.22 4.08 4.79
$\times$ DeWave 2.35 1.25 1.16 1.17 6.24 3.88 3.94 4.28
$\checkmark$ DeWave+MAML [8] 2.08 1.25 1.16 1.17 6.24 3.88 3.94 4.28
Figure 4: The cross-subjects performance variance without calibration
In Table 2, we compare with MAML by reporting the average performance drop ratio between within-subject and cross-subject translation metrics on 18 human subjects on both eye-fixation sliced features and raw EEG waves. We compare the DeWave with the baseline under both direct testing (without Calib data) and with MAML (with Calib data). The DeWave model shows superior performance in both settings. To further illustrate the performance variance on different subjects, we train the model by only using the data from subject YAG and test the metrics on all other subjects. The results are illustrated in Figure 4, where the radar chart denotes the performance is stable across different subjects. Please refer to Appendix F for more detailed results.
@@ 4.5 Generated Samples
Table 3: Translation results on the unseen EEG waves, where bold denotes a correct match between ground truth and our prediction. Underline denotes a fuzzy match with similar semantic meanings.
@@ 5 Conclusion
This paper presents DeWave, a novel framework for the recently proposed open-vocabulary EEG-to-Text translation task [47], introducing the concept of discrete codex encoding. This approach brings about a $1\%$-$10\%$ enhancement in corpus text relevancy metrics, such as BLEU and ROUGE. DeWave also expands the task setting to decode raw EEG waves without the assistance of eye fixation markers, offering a viable solution for the first time to enable direct translation. Despite these advancements, the quality of brain decoding results remains substantially lower than that of language translation and speech recognition. The translation of thoughts directly from the brain is a valuable yet challenging endeavor that warrants significant continued efforts. Given the rapid advancement of Large Language Models, similar encoding methods that bridge brain activity with natural language deserve increased attention.
Limitations  Despite DeWave’s enhancements in EEG-to-Text translation using a discrete codex and raw wave encoding, its accuracy remains inferior compared to traditional language-to-language translations. Additionally, its experiments, restricted to public neural reading data, don’t fully align with the "silent speech" concept of direct thought translation from human brains.
Broader Impacts  DeWave’s introduction of discrete encoding, similar to word embedding, could be a good step in narrowing the gap between brain waves and advanced language models.
@@
Figure 6: Overall illustration of translate EEG waves into text through quantised encoding.
@@ Appendix A Dataset
DeWave utilize both ZuCo [14, 16] 1.0 [14] and 2.0 [16] for our experiments. The ZuCo dataset simultaneously recorded the text and EEG waves corpus during Normal Reading (NR) and Task-Specific Reading (TSR) tasks, where the reading material is from movie reviews or Wikipedia articles. The EEG features are collected with a 128-channel system under the sampling rate 500Hz through a frequency band filter from 0.1Hz to 100Hz. However, after the noise canceling process, only 105 channels [14] are used for translation.
For word-level EEG feature translation, the ZuCo dataset provides eye-fixation data associated with each word while reading. Similar to  [47], we only slice the part of the EEG wave according to the eye fixation. For words with fixation more than once, we concatenate all the fragments for processing. For word-level EEG feature processing, we calculate two statistical results of four frequency band filters, Theta band (5-7Hz), the Alpha band (8-13Hz), the Beta band (12-30Hz), and Gamma band (30Hz-) [25] to get the statistic frequency features of each fragment. For each word, the feature size is $105\times 4\times 2=840$. For raw EEG waves, the signal is normalized into value range $0\sim 1$ for decoding. The reading task’s data are divided into the train ($80\%$), development ($10\%$), and test ($10\%$) respectively by 10874, 1387, and 1387 unique sentences with no intersections.

### v3

@@ Abstract
The translation of brain dynamics into natural language is pivotal for brain-computer interfaces (BCIs). With the swift advancement of large language models, such as ChatGPT, the need to bridge the gap between the brain and languages becomes increasingly pressing. Current methods, however, require eye-tracking fixations or event markers to segment brain dynamics into word-level features, which can restrict the practical application of these systems. To tackle these issues, we introduce a novel framework, DeWave, that integrates discrete encoding sequences into open-vocabulary EEG-to-text translation tasks. DeWave uses a quantized variational encoder to derive discrete codex encoding and align it with pre-trained language models. This discrete codex representation brings forth two advantages: 1) it realizes translation on raw waves without marker by introducing text-EEG contrastive alignment training, and 2) it alleviates the interference caused by individual differences in EEG waves through an invariant discrete codex with or without markers. Our model surpasses the previous baseline (40.1 and 31.7) by $3.06\%$ and $6.34\%$, respectively, achieving 41.35 BLEU-1 and 33.71 Rouge-F on the ZuCo Dataset. This work is the first to facilitate the translation of entire EEG signal periods without word-level order markers (e.g., eye fixations), scoring 20.5 BLEU-1 and 29.5 Rouge-1 on the ZuCo Dataset.
Ground Truth Bob attended the University of Texas at Austin where he graduated, Phi Beta Kappa with a Bachelor’s degree in Latin American Studies in 1973, taking only two and a half years to complete his work, and obtaining generally excellent grades. Predict was the University of California at Austin in where he studied in Beta Kappa in a degree of degree in history American Studies in 1975. and a one classes a half years to complete the degree. and was a excellent grades.
Figure 1: Overall illustration of translating EEG waves into text  [^2] This visualization still keeps teacher-forcing evaluation for fair comparison with previous methods.through quantised encoding.
@@ 1 Introduction
Decoding brain states into comprehensible representations has long been a focal point of research [19, 36, 8, 52]. Electroencephalogram (EEG) signals are particularly favored by researchers due to their non-invasive nature and ease of recording [29, 44]. Traditional EEG decoding techniques largely focus on classifying brain states into restricted categories like Motor Imaginary (MI) [35, 7], Emotion [18, 45], Robotic Control [42], and Gaming [26, 22]. However, these labels, bound to specific tasks, are insufficient for broad-based brain-computer communication. Consequently, there has been a surge of interest in brain-to-text (speech) translation in recent years. As the current trend leans towards large language models (LLM) [4, 11, 28] showcasing increasingly generalized intelligence capabilities, it becomes crucial to delve into ways of bridging the gap between brain signals and natural language representation. However, this area remains under-explored.
@@ 2 Related Works
The key to decoding natural language from EEG signals is good representations. Existing work for EEG-to-text representation can be categorized into hand-crafted representations and deep-learning representations. Earlier works make use of traditional methods to obtain hand-crafted representations from a sequence of EEG signals. A continuous feature representation is extracted using methods such as statistical features [51], correlation coefficient matrix from fast Fourier transform (FFT) components [41, 49, 48], wavelet transform features [17, 30, 52], and Mel-frequency cepstral coefficients (MFCCs) [5]. These hand-crafted representations can be used to establish a mapping between EEG segments and words using distance-based methods such as linear discriminative analysis (LDA) or Support Vector Machine (SVM).
@@ 3.3 EEG Vectorization
Word-Level EEG Features With Event Markers:  The EEG waves are first sliced into fragments according to the eye-tracking fixation of word sequences given in the annotation. Similar to [46], we calculate the statistical result of four frequency band filters, Theta band (5-7Hz), the Alpha band (8-13Hz), the Beta band (12-30Hz), and Gamma band (30Hz-) [25] to get the statistic frequency features of each fragment. It is noted that although different fragments may have different EEG window sizes, the statistical results are the same (embedding size 840). A multi-head transformer layer is applied to project the embedding into feature sequences with latent size 512.
@@
For structure-wise, a conformer-based multi-layer encoder with specially designed hyperparameters is employed. The one-dimensional convolution layer processes the EEG waves to generate the embedding sequence 444perception field is roughly 200ms with overlap 100ms for each embedding, fusing the EEG channels into a unique embedding for each period. We apply bi-directional transformer attention layers to the sequence to capture temporal relations.
@@
To obtain a semantically coherent codex, we introduce a cross-modality contrastive learning approach distinct from prevalent methods. Unlike CLIP [33, 21] that contrasts CLS embeddings between sample pairs within a mini-batch, our approach operates within a single EEG-text pair. We contrast the EEG-tokenized codex embeddings sequence $z_{q}$ with the text embeddings sequence $z_{t}$. Assuming the raw wave feature extractors can produce a token sequence in an organized chronological order, we treat the diagonal EEG codex and text word2vec encoding pairs as positive pairs within the sequences. All other pairs are considered negative. The model is trained to minimize the distance between embeddings of positive pairs and maximize that of the negative pairs. For a given EEG-text pair $(i,j)$, the loss is defined in Equation 4
$L_{\rm{contrast}}=-\frac{1}{n}\sum log\left[\frac{\exp(\mathbf{s}_{ii}/\tau)}{\sum_{k=1}^{N}\exp(\mathbf{s}_{ik}/\tau)}\right],\mathbf{s}_{ij}=\mathbf{z}_{q}(\mathbf{x}^{i})^{T}\mathbf{z}_{t}(j)$ (4)
@@ 4.2 Implementation Details
For word-level EEG features, we use the $56$ tokens each with an $840$ embedding size. For raw EEG waves, we clip or pad the EEG waves up to sample point 5500 with a constant value of zero. A transformer layer with head number 8 and a $1\times 1$ convolutional layer are combined to fuse multiple EEG channels into an embedding sequence with size 512. DeWave uses a codex with size 2048 where each codex latent is an embedding with size 512. The ablation study (Section 4.6) gives a discussion about the codex size. All models are trained on Nvidia V100 and A100 GPUs. For the self-supervised decoding for raw waves, we use a learning rate of 5e-4 and a VQ coefficient of 0.25 for training 35 epochs. For training the codex (stage 1), DeWave uses a learning rate of 5e-4 for 35 epochs. For finetuning the translation (stage 2), DeWave uses a learning rate of 5e-6 for 30 epochs. We use the SGD as the optimizer for training all the models. Due to limited space, refer to Appendix B for more details.
@@ 4.3 Evaluation Metrics
We evaluate translation performance using NLP metrics, BLEU and ROUGE, as shown in Table 1. For word-level EEG features, we compare our results to EEG-to-Text [46], maintaining a consistent language model for fairness. In the absence of methods for raw EEG waves, we establish a baseline ${}^{\dagger}$ by segmenting the entire EEG waves into a sequence embedding using a 200ms time window with a 100ms overlap. We adapt Wave2Vec, originally developed for speech recognition, to brain waves and compare it with our approach, DeWave. Furthermore, we adapt unsupervised raw EEG waves classification methods BENDR [10] and SCL [9], using SSL pre-training and feature extraction for comparison, underscoring the impact of discrete encoding.
Table 1: Evaluation metrics of EEG-to-Text translation under both word-level features input and raw waves input, where +Contrastive denotes boost encoder with $L_{contrast}$ mentioned in Sec. 3.3. For a fair comparison, these results keep the same teacher-forcing evaluation setting as EEG-to-Text [46].
@@
Raw Waves:   DeWave represents an unprecedented endeavor in translating raw EEG waves directly, obviating the need for event markers. This is achieved through the application of a learned Wave2Vec model, which converts waves into discrete codex tokens. When benchmarked against the baseline method - which merely slices the wave according to a time window - DeWave exhibits marked improvement. It attains BLEU-$1,2,3,4$ scores of $20.51,10.18,5.16$, and $2.56$, respectively, surpassing the baseline by margins of $7.44~{}(+56.92\%)$, $4.44~{}(+76.1\%)$, $2.61~{}(+102.35\%)$, and $1.42~{}(+129.09\%)$. This underscores the superiority of a learned discrete embedding representation over the rudimentary time-window baseline. In comparison with cutting-edge self-supervised learning (SSL) methods, DeWave consistently outperforms them, including contrastive learning (CL)-based methods such as SCL [9], the original Wave2Vec [2], and BENDR [10].
@@ 4.4 Cross-Subject Performance
Cross-subject performance is of vital importance for practical usage. To further illustrate the performance variance on different subjects, we train the model by only using the data from subject YAG and test the metrics on all other subjects. The results are illustrated in Figure 4, where the radar chart denotes the performance is stable across different subjects. Please refer to Appendix G for more detailed results.
Figure 4: The cross-subjects performance on ZuCo dataset.
@@ 4.5 Generated Samples
Table 2: Translation results on the unseen EEG waves, where bold denotes a correct match between ground truth and our prediction. Underline denotes a fuzzy match with similar semantic meanings. For a fair comparison, these results keep the same teacher-forcing evaluation setting as EEG-to-Text [46]. This means the decoding process eliminates accumulated errors and just predicts the current token with the GT token from the last step.
@@ 4.6 Ablation Study
5 Limitations
Despite DeWave’s enhancements in EEG-to-Text translation using a discrete codex and raw wave encoding, its accuracy remains far from real-life scenarios compared to traditional language-to-language translations. Also, to keep a fair comparison with EEG-to-Text, this paper uses a teacher-forcing setting when evaluating. This setting eliminates accumulation error and turns the sequence decoding task into a word-level classification task given the ground truth token from the previous step. This setting is much easier yet we think it is still valuable as it could suggest feature extraction quality while keeping a fair comparison.
Additionally, the experiments in this paper are restricted to public neural reading data, not fully align with the "silent speech" concept of direct thought translation from human brains. Instead, the current ZuCo dataset is collected by giving people reading stimuli. In that case, this is still “passively" trying to reconstruct the stimuli signal. This paper focuses on introducing Wav2Vec formation of feature extraction on raw EEG waves and introduces discrete codex as learnable representations for the EEG-to-Text translation domain. The scientific problem of which is a better way of doing the “silent speech" task remains unexplored for future works.
@@ 6 Conclusion
This paper presents DeWave, a framework for the recently proposed open-vocabulary EEG-to-Text translation task [46], introducing the concept of discrete codex encoding. This approach brings enhancement in corpus text relevancy metrics, such as BLEU and ROUGE. DeWave also expands the task setting to decode raw EEG waves without the assistance of eye fixation markers, offering a viable solution for the first time to enable direct translation. Despite these advancements, the quality of brain decoding results remains substantially low and still remains teacher forcing setting for fair comparison. The translation of thoughts directly from the brain is a valuable yet challenging endeavor that warrants significant continued efforts. Given the rapid advancement of Large Language Models, similar encoding methods that bridge brain activity with natural language deserve increased attention. In our future work, we will explore more reasonable settings that remove teacher forcing for both training and testing. Also include neural-feedback into this domain to increase the scientific nature of this domain.
Acknowledgement
This work was supported in part by the Australian Research Council (ARC) under discovery grant DP210101093 and DP220100803, and the GrapheneX-UTS Human-Centric AI Centre sponsored by GrapheneX (2023-2031).
@@ Acknowledgement
Figure 6: Overall illustration of translating EEG waves into text through quantised encoding.
@@ Appendix A Dataset
ZuCo stands for Zurich Cognitive Language Processing Corpus (ZuCo), a dataset that includes both raw and preprocessed eye-tracking and electroencephalography (EEG) data. The data is collected by having human subjects read given text corpora while simultaneously recording both their eye-tracking signals and EEG waves. The recording is done using the Biosemi-128 system, which, after denoising, provides 105 out of 128 channels for downstream tasks. The dataset comprises two versions: ZuCo 1.0, collected from 12 subjects, and ZuCo 2.0, collected from 18 subjects [14, 16].
The text corpora within the ZuCo dataset are sourced from a diverse set of textual genres, including 1) Wikipedia articles, 2) movie reviews, and 3) the BNC (British National Corpus). This diversity ensures a wide variety of syntactic structures and word frequencies. The dataset records data during two tasks: Normal Reading (NR) and Task-Specific Reading (TSR). In our experiments, DeWave utilizes both ZuCo 1.0 [14] and 2.0 [16]. The EEG features are captured using a 128-channel system with a sampling rate of 500Hz, filtered through a frequency band ranging from 0.1Hz to 100Hz. After noise canceling, only 105 channels are deemed suitable for translation [14].
For word-level EEG feature translation, eye-fixation data associated with each word during reading is available in the ZuCo dataset. Following the approach similar to  [46], we extract segments of the EEG wave according to eye fixations. Words fixated upon multiple times have their EEG fragments concatenated for processing. To process these word-level EEG features, we compute statistical results across four frequency band filters: the Theta band (5-7Hz), the Alpha band (8-13Hz), the Beta band (12-30Hz), and the Gamma band (30Hz and above) [25]. Consequently, the feature size for each word totals $105\times 4\times 2=840$. For raw EEG waves, the signals are normalized to a range between 0 and 1 for decoding.
The dataset is split into training (80%), development (10%), and testing (10%) sets, comprising 10,874, 1,387, and 1,387 unique sentences, respectively, with no overlap. We further conducted a statistical analysis on the sentences extracted from the dataset, details of which are reported below.
Table 4: Statistical analysis of sentences from the ZuCo dataset.
Feature ZuCo 1.0 Natural Reading ZuCo 2.0 Natural Reading
Sentences 300 390
Sent. length 21.3 $\pm$ 10.6 19.6 $\pm$ 8.8
Total words 6386 6828
Word length 6.7 4.9